OpenStudy (anonymous):

PLEASE HELP, I WILL FAN IF YOU GET THE ANSWER RIGHT!!!! Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (crashonce):

ok the equation of a line is (y-y1)=m(x=x1) now try

OpenStudy (anonymous):

So y - -3 (which makes it +) y + 3 = -6(x +2)

OpenStudy (anonymous):

I am not sure if that's the formula I am supposed to use?

OpenStudy (crashonce):

y1 = -3, x1=-2

OpenStudy (crashonce):

m=-6 now substitute

OpenStudy (texaschic101):

or just use y = mx + b with the slope (m) being -6 and use the points given (-2,-3)...x = -2 and y = -3.....and sub that back into the formula and solve for b

OpenStudy (anonymous):

-3 = -6(-2)

OpenStudy (anonymous):

so 12?

OpenStudy (texaschic101):

-3 = -6(-2) + b.....now solve for b

OpenStudy (crashonce):

y=y1 so it will be positive

OpenStudy (anonymous):

how do I solve for b?

OpenStudy (crashonce):

y-y1=m(x-x1) y-(-3)=-6(x-(-2))

OpenStudy (crashonce):

simplify

OpenStudy (anonymous):

y + 3 = -6(x + 2)

OpenStudy (anonymous):

but how do I solve it from there?

OpenStudy (crashonce):

expand the RHS and simplify

OpenStudy (texaschic101):

your probably getting confused here....you are seeing two different ways to do this problem...I will be quiet and let CrashOnce finish explaining

OpenStudy (crashonce):

ok basically expand the RHS and make it look like the answers so have just 'y' on the LHS

OpenStudy (anonymous):

Yea it is confusing, thank you :)

OpenStudy (anonymous):

I haven't learned about RHS and LHS so it's confusing to be explained that way

OpenStudy (crashonce):

RHS=right hand side LHS= left hand side lol

OpenStudy (anonymous):

Oh, lol thanks

OpenStudy (crashonce):

no worries

OpenStudy (anonymous):

what I don't understand is the answer choices

OpenStudy (anonymous):

I'm not sure where they fit in?

OpenStudy (crashonce):

y-(-3)=-6(x-(-2)) y+3=-6(x+2) y+3=-6x-12 therefore y=-6x-12-3 y=-6x-15

OpenStudy (crashonce):

i shouldnt be posting answers but does that make more sense?

OpenStudy (anonymous):

Ohhhhhh I forgot to multiply -6 by x and 2 that were in the parenthesis... I have been doing that a lot

OpenStudy (crashonce):

lol take care with brackets

OpenStudy (crashonce):

solve the brackets first

OpenStudy (anonymous):

Yes! Thank you! sorry for taking so long to understand :)

OpenStudy (texaschic101):

I am not trying to confuse you and mess with your brain..but can I show you the way I do it....I find it a little bit easier. Are you interested ?

OpenStudy (anonymous):

Yes of course! I would like to know other ways as well

OpenStudy (crashonce):

ill 'retreat' now thanks c:

OpenStudy (anonymous):

haha okay, thank you! @CrashOnce

OpenStudy (texaschic101):

the formula we will use is y = mx + b the slope is always in the m position. The points (-2,-3)...x = -2 and y = -3 now we sub into the formula y = mx + b -3 = -6(-2) + b -3 = 12 + b -3 - 12 = b -15 = b so we now have : y = -6x + (-15) or y = -6x - 15

OpenStudy (anonymous):

That makes a lot of sense too! I was always going back to the y = mx + b formula, I just wasn't sure what to do with b...Thank you for takin your time to respond! All the help is much appreciated! :) @texaschic101

OpenStudy (texaschic101):

your very welcome :) you can also do it the way CrashOnce did it

OpenStudy (anonymous):

Okay, thank you! :)